EP018A

PI: Paragi
Title: 3C287

Session: Feb 1997 (session 1)
Frequency: 6 cm
Mode: B
Obs time (UT): EU 20:30(23/02)-07:00(24/02)
Stations: EbJbMcNtOnShSmTrUrWb
Processor: Bonn
EU Mk3 Tapes/St.: 2

Comment


To give feedback on this experiment, click here

Experiment Feedback

Medicina (T. Venturi at Mon Feb 24 02:13:55 EST 1997) Success
The experiment ran smoothly. No problems encountered.

Effelsberg (dgraham at Mon Feb 24 03:33:00 EST 1997) Success
OK

Onsala (Antonis Polatidis at Mon Feb 24 05:08:31 EST 1997) Success
Everything looks OK

Onsala (Antonis Polatidis at Mon Feb 24 05:08:37 EST 1997) Success
Everything looks OK

Westerbork (Tony Foley at Mon Feb 24 11:02:42 EST 1997) Minor failures
Restart at 21:28 UT may give phase jump. Tape stopped at 22:58UT because it claimed it had a vacuum loss (NOT true!) and later (23:12UT) it ran off the reel. Restarted by operator at 02:20.

Shanghai (Huang Xinyong at Wed Mar 5 05:01:16 EST 1997) Success
Clock offset = 0.27 - 0.29 msec. No problems, weather cloudy.

Jodrell (Alastair Gunn at Thu Mar 6 03:32:49 EST 1997) Minor failures
Telescope parked at 23:10 for rest of experiment due to high winds. Power failure occurred at about 07:00 UT, clocks reset. Tapes RRL00068,0294 and ASI00031,F078 sent to Bonn via Woodcroft on 24/02/97.